# Gauging or extending bulk and boundary conformal field theories: Application to bulk and domain wall problem in topological matter and their descriptions by (mock) modular covariant

### v2

@@ III.2 Anomalous theory and massive flow: spontaneous symmetry breaking versus breaking of extended symmetry
Because the theory is anomalous, the parity index $p$ carries charge. Hence the original bosonic smeared Cardy state $|i,p\rangle\rangle$ has different charges depending both on $i$ and $p$. This is a modified LSM-type argument using smeared Cardy states. One can see related arguments in [245, 246, 195, 247, 248, 132].
Figure 6: Existence of disordered phase and its implication in an anomalous theory.
In the parafermionic models, the disordered phase is more nontrivial from this view. For simplicity, let us concentrate our attention on the three-state Potts model. If one admits the conjecture by Cardy for this case, one needs to implement $Z_{N}$ invariant boundary states. Moreover, the resultant Cardy state should be related to the SSB phase under the Kramers-Wannier transformation. Hence by applying the analogy to the Ising model, one has to introduce the free boundary condition $|\sigma\rangle\rangle$ as $Z_{3}$ invariant boundary state. In literature, the corresponding operator $\sigma$ is a field in the $Z_{2}$ orbifolded theory of the three-state Potts model known as minimal model $M(6,7)$[249] and this is outside of the primary fields of the original three-state Potts CFT. Whereas the original theory is an anomalous $Z_{3}$ model, the orbifolded theory is an anomaly-free $Z_{2}$ theory. In other words, the extended models appear as a disordered phase. However, in general, the extended object should break extended symmetry such as W-symmetry or Kac-Moody Lie group symmetry. These extended boundary conditions are called symmetry-breaking branes or symmetry-breaking boundary conditions in high energy physics [250, 251, 252, 253, 254, 255] and new boundary conditions in condensed matter or statistical mechanical physics[249, 256, 257] (we note the textbook [156] for reference). Based on this argument, one can understand the appearance of a disordered gapped phase in an anomalous system as a signal of the existence of an extended model (Fig. 6). Hence such anomalous theory can be classified by the extended model (or anomalous theory should be considered as a consequence of the orbifolding of anomaly-free model).
@@ IV Massless flow and its implications in topological order
In the previous section, we concentrated on the RG flow from CFTs to the gapped phase in $1+1$ dimensions. Under CFT/TQFT, this is equivalent to the bulk property of TOs in $2+1$ dimensions. However, in RGs from a CFT, there exists another situation where the system flows to other CFTs, called massless RG flow[101]. In this section, we revisit the problem by studying the tensor product of the CFTs connected by RG flow under preserving $Z_{N}$ symmetry. The construction of modular partition functions of coupled topologically ordered systems is proposed from a modern view and they contain a new series of modular partition functions that correspond to FQHE with nonchiral anyons. We show two types of theories by the different anomaly classification mechanisms, anomaly cancellation for the usual coupled (or layered) model and anomaly matching for the nonchiral theory. The Halperin states[142] or the FQHE in HLL are typical examples of the former and latter cases respectively. Moreover, by using the folding trick, we demonstrate that even the known partition can be interpreted as a nontrivial domain wall changing the chirality of anyons.
@@
The above partition function corresponds to $Z_{N}$ extension of the SFC $\mathbf{B}_{\text{UV}}\otimes\mathbf{B}_{\text{IR}}$ induced by the $Z_{N}$ group generator $J_{\text{tot}}$. Hence we denote the fusion algebra of this extended theory as $\mathbf{F}_{\text{UV}\times\text{IR}}$. We also note that this algebra is included in $\mathbf{F}_{\text{UV}}\otimes\mathbf{F}_{\text{IR}}$ but they are not equivalent. Depending on the class of the theories or the form of $J_{\text{tot}}$, the character $\Xi$ becomes chiral or nonchiral and this corresponds to the non-chiral anyon in [21]. The total charge $Q_{J_{\text{tot}}}$ is the generalized $Z_{N}$ charge (or anomaly) defined by the conformal spin (or by conformal dimensions indirectly) of the operators. For simplicity, we concentrate our attention on the unitary models, but one can generalize the arguments to nonunitary models, by replacing the definition of charge as monodromy charge[162] (For the readers interested in the application of nonunitary CFTs to TOs, we note works on gapless FQHEs [258, 259, 260, 261, 262, 263]).
@@
  • ($N$ integer) $s_{J_{\text{tot}}^{k}}=$integer for all $k$ and $N$,
  • ($N$ even case) $s_{J_{\text{tot}}^{k}}=$ integer for $k$: even and $s_{J_{\text{tot}}^{k}}=$ half-integer for $k$: odd,
@@ IV.1 Massless flow as coupled topological orders
Because $Z_{\text{UV}}$ and $Z_{\text{IR}}$ are modular invariant, one can implement the product $Z_{\text{UV}}Z_{\text{IR}}$ as a modular invariant. However, the partition function $Z_{\text{UV}}Z_{\text{IR}}$ is a trivial theory and this theory itself does not produce a nontrivial conformal interface. Here we introduce one type of anomaly freeness of the product theory,
  • ($N$: Integer) $h_{J^{k}_{\text{UV}}J^{k}_{\text{IR}}}=$ integer for all $k$ and $N$
  • ($N$: Even integer) $h_{J^{k}_{\text{UV}}J^{k}_{\text{IR}}}$= integer for $k$: even and $h_{J^{k}_{\text{UV}}J^{k}_{\text{IR}}}$= half-integer for $k$: odd
@@
where the total charge of each sector $Q_{\text{UV}\times\text{IR}}=Q_{\text{UV}}+Q_{\text{IR}}$ becomes zero. Similarly one can introduce the twisted partition function as
@@
  • ($N$ integer) $s_{(j\overline{j})^{k}}=$integer for all $k$ and $N$,
  • ($N$ even case) $s_{{(j\overline{j})}^{k}}=$ integer for $k$: even and $s_{(j\overline{j})^{k}}=$ half-integer for $k$: odd,
@@
for an anomaly-free UV theory where we have taken the summation for the $i$ and $a$ indices in both UV and IR theories. By taking the total charge $Q_{\text{UV}\times\text{IR}}$ to 00, one obtains a modular invariant.
@@
with $Z_{N}$ electron operator $j\overline{j}$. The single-valuedness of the wavefunctions is satisfied by the anomaly-freeness of the operator $j\overline{j}$. Hence the resultant wavefunctions in plane geometry can contain both contributions from holomorphic and and antiholomorphic variables. To distinguish this theory from a usual CCFT appearing as edge modes of topological order, we call this as one edge CFT (OECFT). As far as we know, a similar mixing of the holomorphic and antiholomorphic pairs appears in the study of FQHEs with nonchiral anyon, such as antiPfaffian and HLL. Hence, we propose that FQHE with nonchiral anyon is described by OECFT corresponding to anomaly-matched RG flow. Similar observations can be seen in [11], and our proposal can be considered as a generalization of their arguments  [^3] We thank Justin Kaidi for related discussions.
Figure 7: OECFT constructed from the product of $Z_{N}$ extended theories.
@@ IV.2 Massless flow as the interface of topological orders
In the previous subsection, we studied the products of theories connected by the RG domain wall. By applying the folding trick[117] to this coupled model, it is possible to interpret the set of partition functions as a mapping of theories. Hence there exists following mapping $D_{\text{RG}}$ corresponding to the modular invariant or the charge 00 sector of the product of the theories $\{\mathbf{F}_{\text{UV}}\otimes\mathbf{F}_{\text{IR}}\}_{Q_{\text{tot}}=0}$,
@@
where $\mathbf{F}$ is the $Z_{N}$ extension of SFC corresponding to $Z_{N}$ extended bulk CFT. We note that it is possible to implement the domain walls that do not preserve the anomaly by applying the folding trick to the coupled model with the charged partition functions. It may be reasonable to call them $Z_{N}$ charged domain walls analogous to the domain walls studied in astrophysics[282, 283, 284, 285, 286, 287]. We concentrate our attention on $Q_{\text{tot}}=0$ for simplicity, but by attaching charged particle (analogous to quark) to the domain wall $Q_{\text{tot}}=0$ or studying the structure of modular covariant $Q_{\text{tot}}\neq 0$, one can implement the charged domain wall. We also remark that one can generalize the setting to multi-domain walls and shrink them freely (Fig. 8). Hence by this construction, one can obtain the possible set of domain wall particles systematically. We also note recent studies on domain wall in condensed matter [115, 288, 116, 289, 290, 291]. In principle, one can apply our analysis to the corresponding quantum multi-junction problems (or Kondo problem)[292, 293, 294, 295, 296, 297, 298, 299, 300, 301]. For the readers interested in the theoretical analysis of junction problem, we note two reviews[302, 303].
Figure 8: Multiple domain walls and the consequence of shrinking. One can map anyons in one system to another, and change the shape of the system consistently. By changing the geometry of the settings, one can apply our analysis to multi-junction problems (To obtain analytical data, more respective calculations are necessary). Similar ideas can be seen commonly in literature[300, 115, 116, 289, 290].
Roughly speaking, one can identify Eq. (80) providing a mapping relation between two topological orders in a cylinder geometry. Because the Witt equivalence of existing MTCs is too strong to apply for the analysis of the RG domain wall, we have introduced $Z_{N}$ extension of SFCs and defined its anomaly-free conditions as an equivalence relation. We also remind that the fundamental point connecting the phenomenology of interface and that of the product theory is the folding trick[117]. Hence the role of chiral and antichiral parts of either $UV$ or $IR$ theory should be swapped. Roughly speaking, one can understand each object appearing in the set of modular partition functions on the basis of $\{\mathbf{F}_{\text{UV}}\otimes\mathbf{F}_{\text{IR}}\}_{Q_{\text{tot}}=0}$ as the following mapping,
@@
where $\overline{\alpha}$ is the complex conjugation of $\alpha$ and we droped the parity indeces for simplicity. However, this only implies the condition for the charge and parity, and more detailed data are necessary for pratical calculations. Hence the following discussion, we assume that the construction of the RG domain wall has been established by more respective techniques as in [102, 119, 103, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 15] and we concentrate our attention on the properties obtained from the charge and parity structures. Surprisingly, only from these structures, one can obtain nontrivial domain walls changing the chirality of anyons.
@@
The right arrow $\Rightarrow$ is obtained by bulk semionization or sandwich construction. To check consistency between this formalism and the domain wall Verlinde formula in [116] (or similar generalization [304]) is an interesting research direction. For simplicity, we study the models with anomalous UV and IR theories, but the same phenomenolgy should hold for the models with the anomaly free UV and IR theories.
@@
where $(i)$ is the label of two systems and $\pm$ is labeling chiral and anithiral components. Hence by applying the folding trick[117], the two theories are Witt equivalent when $J_{\text{tot}}=J^{(1)}\overline{J}^{(2)}$ satisfy the anomaly-free condition.
@@
[…]
@@ V Conclusion
Figure 10: Respective benefit of the bosonic representations and group extended representations. In a bosonic model, a correspondence between operators constructing correlation functions, quantum states, and (generalized) symmetry holds. Hence one can achieve various calculations, by using vertex-operator-algebra or category theory and more. In a group-extended model analogous to a hadron model (or fractional supersymmetric model), the theory contains information on anomaly classifications determined by the conformal dimensions of the simple current. Moreover, one can attach gauge field easily. However, because the zero modes contain the nontrivial entanglement, calculations of the correlation functions become difficult.
@@
As a concluding remark, we note the respective benefit of the bosonic and fermionic (or group extended) representations. Bosonic models have been extensively studied because this representation usually provides systematic calculations of correlation functions and related physical observables. In a modern understanding, the correspondence between quantum states, correlators, and (generalized) symmetry seems to be a fundamental source of this benefit. On the other hand, the group-extended model corresponds to a particle (or quark-hadron) picture. Hence one can attach gauge fields and apply the ideas of condensations more easily (Fig. 10). However, as we have observed in this work, the treatment of the zero modes and their entanglement structure requires careful treatment. We also remark on the difficulty of attaching gauge fields to bosonic systems[323, 324, 325, 326, 327, 328, 329, 330, 331, 332]. (It may be surprising for the readers, but the naive coupling of bosonic fields and $U(1)$ gauge field cannot reproduce the LSM theorem and the spectrum of twisted models[333, 334].) Because of the theoretical difficulties of the gauging operations, studies on the corresponding lattice or microscopic models are fundamentally important respectively. Inversely, the field-theoretical understanding of such models is also important for more general understanding of quantum phases.

### v3

@@ III.2 Anomalous theory and massive flow: spontaneous symmetry breaking versus breaking of extended symmetry
Because the theory is anomalous, the parity index $p$ carries charge. Hence, the original bosonic smeared Cardy state $|i,p\rangle\rangle$ has different charges depending on both $i$ and $p$. This is a modified LSM-type argument using smeared Cardy states. One can see related arguments in [249, 250, 197, 251, 252, 133].
Figure 6: Existence of disordered phase and its implication in an anomalous theory. If an anomalous $Z_{N}$ symmetric system flows to a nondegenerate gapped phase labelled by $a$, this indicates the existence of symmetry-breaking boundary condition in the original CFT. In other words, the information of massive RGs will provide nontrivial data of new or unfamiliar BCFTs.
In the parafermionic models, the disordered phase is more nontrivial from this view. For simplicity, let us concentrate our attention on the three-state Potts model. If one admits the conjecture by Cardy for this case, one needs to implement $Z_{N}$ invariant boundary states. Moreover, the resultant Cardy state should be related to the SSB phase under the Kramers-Wannier transformation. Hence, by applying the analogy to the Ising model, one has to introduce the free boundary condition $|\sigma\rangle\rangle$ as a $Z_{3}$ invariant boundary state. In literature, the corresponding operator $\sigma$ is a field in the $Z_{2}$ orbifolded theory of the three-state Potts model known as the minimal model M(6,7)M(6,7)[253], and this is outside of the primary fields of the original three-state Potts CFT. Whereas the original theory is an anomalous $Z_{3}$ model, the orbifolded theory is an anomaly-free $Z_{2}$ theory. In other words, the extended models appear as a disordered phase. However, in general, the extended object should break extended symmetry such as W-symmetry or Kac-Moody Lie group symmetry. These extended boundary conditions are called symmetry-breaking branes or symmetry-breaking boundary conditions in high energy physics [254, 255, 256, 257, 258, 259] and new boundary conditions in condensed matter or statistical mechanical physics[253, 260, 261] (we note the textbook [157] for reference). Based on this argument, one can understand the appearance of a disordered gapped phase in an anomalous system as a signal of the existence of an extended model (Fig. 6). Hence such an anomalous theory can be classified by the extended model (or anomalous theory should be considered as a consequence of the orbifolding of an anomaly-free model).
@@ IV Massless flow and its implications in topological order
In the previous section, we concentrated on the RG flow from CFTs to the gapped phase in $1+1$ dimensions. Under CFT/TQFT, this is equivalent to the bulk property of TOs in $2+1$ dimensions. However, in RGs from a CFT, there exists another situation where the system flows to other CFTs, called massless RG flow[101]. In this section, we revisit the problem by studying the tensor product of the CFTs connected by RG flow while preserving $Z_{N}$ symmetry. The construction of modular partition functions of coupled topologically ordered systems is proposed from a modern view and they contain a new series of modular partition functions that correspond to FQHE with nonchiral anyons. We show two types of theories by the different anomaly classification mechanisms, anomaly cancellation for the usual coupled (or layered) model and anomaly matching for the nonchiral theory. The Halperin states[143] or the FQHE in HLL are typical examples of the former and latter cases respectively. Moreover, by using the folding trick, we demonstrate that even the known partition can be interpreted as a nontrivial domain wall changing the chirality of anyons.
@@
The above partition function corresponds to $Z_{N}$ extension of the SFC $\mathbf{B}_{\text{UV}}\otimes\mathbf{B}_{\text{IR}}$ induced by the $Z_{N}$ group generator $J_{\text{tot}}$. Hence we denote the fusion algebra of this extended theory as $\mathbf{F}_{\text{UV}\times\text{IR}}$ (Recently the structure corresponding to $\mathbf{F}_{\text{UV}\times\text{IR}}$ has been studied in [262], and this work contains useful calculations of current operators. However, the corresponding coset representations have already been studied in [104] and the resultant partition function in the present work is more general). In the pioneering work [104], it has been conjectured that the BCFT of the extended theory $\mathbf{F}_{\text{UV}\times\text{IR}}$ in the present work produces the information of the massless RG between the original $\mathbf{B}_{\text{UV}}$ and $\mathbf{B}_{\text{UV}}$ theories (one can see the phenomenology in the figure 2 of [104] in a concise way). We also remind that the appearance of extended model for the classification of BCFTs is common in the studies of symmetry breaking boundary conditions[253, 157]. We also note that this algebra $\mathbf{F}_{\text{UV}\times\text{IR}}$ is included in $\mathbf{F}_{\text{UV}}\otimes\mathbf{F}_{\text{IR}}$ but they are not equivalent. Depending on the class of the theories or the form of $J_{\text{tot}}$, the character $\Xi$ becomes chiral or nonchiral and this corresponds to the non-chiral anyon in [21]. The total charge $Q_{J_{\text{tot}}}$ is the generalized $Z_{N}$ charge (or anomaly) defined by the conformal spin (or by conformal dimensions indirectly) of the operators. For simplicity, we concentrate our attention on the unitary models, but one can generalize the arguments to nonunitary models, by replacing the definition of charge as monodromy charge[163] (For the readers interested in the application of nonunitary CFTs to TOs, we note works on gapless FQHEs [263, 264, 265, 266, 267, 268]).
@@
  • (NN integer) $s_{J_{\text{tot}}^{k}}=$integer for all $k$ and $N$,
  • (NN even case) $s_{J_{\text{tot}}^{k}}=$ integer for $k$: even and $s_{J_{\text{tot}}^{k}}=$ half-integer for $k$: odd,
@@ IV.1 Massless flow as coupled topological orders
Because $Z_{\text{UV}}$ and $Z_{\text{IR}}$ are modular invariant, one can implement the product $Z_{\text{UV}}Z_{\text{IR}}$ as a modular invariant. However, the partition function $Z_{\text{UV}}Z_{\text{IR}}$ is a trivial theory because all of the charge sectors in the UV and IR theories mix randomly. Hence, this theory itself does not produce a nontrivial conformal interface with ’t Hooft anomaly matching condition straightforwardly. Here we introduce one type of anomaly freeness of the product theory,
  • (NN: Integer) $h_{J^{k}_{\text{UV}}J^{k}_{\text{IR}}}=$ integer for all $k$ and NN
  • (NN: Even integer) $h_{J^{k}_{\text{UV}}J^{k}_{\text{IR}}}$= integer for $k$: even and $h_{J^{k}_{\text{UV}}J^{k}_{\text{IR}}}$= half-integer for $k$: odd
@@
where the total charge of each sector $Q_{\text{UV}\times\text{IR}}=Q_{\text{UV}}+Q_{\text{IR}}$ becomes zero. Hence, this is more convenient than the original $Z_{\text{UV}}Z_{\text{IR}}$ when studying symmetry (or charge) preserving domain wall. Similarly one can introduce the twisted partition function as
@@
  • (NN integer) $s_{(j\overline{j})^{k}}=$integer for all $k$ and $N$,
  • (NN even case) $s_{{(j\overline{j})}^{k}}=$ integer for $k$: even and $s_{(j\overline{j})^{k}}=$ half-integer for $k$: odd,
@@
for an anomaly-free UV theory where we have taken the summation for the $i$ and $a$ indices in both UV and IR theories. By taking the total charge $Q_{\text{UV}\times\text{IR}}$ to 0, one obtains a modular invariant.
@@
with $Z_{N}$ electron operator $j\overline{j}$. The single-valuedness of the wavefunctions is satisfied by the anomaly-freeness of the operator $j\overline{j}$. Hence, the resultant wavefunctions in plane geometry can contain both contributions from holomorphic and antiholomorphic variables. To distinguish this theory from a usual CCFT appearing as edge modes of topological order, we call this one edge CFT (OECFT). As far as we know, a similar mixing of the holomorphic and antiholomorphic pairs appears in the study of FQHEs with nonchiral anyons, such as antiPfaffian and HLL. Hence, we propose that FQHE with nonchiral anyon is described by OECFT corresponding to anomaly-matched RG flow. Similar observations can be seen in [11], and our proposal can be considered as a generalization of their arguments  [^3] We thank Justin Kaidi for related discussions.
Figure 7: OECFT constructed from the product of $Z_{N}$ extended theories. By interpreting the partition function as that of $CFT_{UV}\otimes CFT_{IR}$ and focusing on the chiral-chiral or chiral antichiral structure $\Xi$, one can obtain the pairing structures of UV and IR theories. The chiral-chiral pairs appear in the systems with anomaly cancellation, and the chiral-antichiral pairs appear in the systems with anomaly matching.
@@ IV.2 Massless flow as the interface of topological orders
In the previous subsection, we studied the products of theories connected by the RG domain wall. By applying the folding trick[118] to this coupled model, it is possible to interpret the set of partition functions as a mapping of theories. Hence, there exists following mapping $D_{\text{RG}}$ corresponding to the modular invariant or the charge 0 sector of the product of the theories $\{\mathbf{F}_{\text{UV}}\otimes\mathbf{F}_{\text{IR}}\}_{Q_{\text{tot}}=0}$,
@@
where $\mathbf{F}$ is the $Z_{N}$ extension of SFC corresponding to $Z_{N}$ extended bulk CFT. We note that it is possible to implement the domain walls that do not preserve the anomaly by applying the folding trick to the coupled model with the charged partition functions. It may be reasonable to call them $Z_{N}$ charged domain walls analogous to the domain walls studied in astrophysics[287, 288, 289, 290, 291, 292]. We concentrate our attention on $Q_{\text{tot}}=0$ for simplicity, but by attaching charged particle (analogous to quark) to the domain wall $Q_{\text{tot}}=0$ or studying the structure of modular covariant $Q_{\text{tot}}\neq 0$, one can implement the charged domain wall. We also remark that one can generalize the setting to multi-domain walls and shrink them freely (Fig. 8). Hence, by this construction, one can obtain the possible set of domain wall particles systematically. We also note recent studies on domain wall in condensed matter [116, 293, 117, 294, 295, 296]. In principle, one can apply our analysis to the corresponding quantum multi-junction problems (or Kondo problem)[297, 298, 299, 300, 301, 302, 303, 304, 305, 306]. For the readers interested in the theoretical analysis of junction problem, we note two reviews[307, 308].
Figure 8: Multiple domain walls and the consequence of shrinking. One can map anyons in one system to another (between $CFT^{(1)}$, $CFT^{(0)}$, and $CFT^{(2)}$ in the figure), and change the shape of the system consistently (resulting in $CFT^{(1)}\rightarrow CFT^{(2)}$). By changing the geometry of the settings, one can apply our analysis to multi-junction problems (To obtain analytical data, more respective calculations are necessary). Similar ideas can be seen commonly in literature[305, 116, 117, 294, 295].
Roughly speaking, one can identify Eq. (80) providing a mapping relation between two topological orders in a cylinder geometry. Because the Witt equivalence of existing MTCs is too strong to apply for the analysis of the RG domain wall, we have introduced $Z_{N}$ extension of SFCs and defined its anomaly-free conditions as an equivalence relation. We also remind that the fundamental point connecting the phenomenology of interface and that of the product theory is the folding trick[118]. Hence, the role of chiral and antichiral parts of either $UV$ or $IR$ theory should be swapped. Roughly speaking, one can understand each object appearing in the set of modular partition functions on the basis of $\{\mathbf{F}_{\text{UV}}\otimes\mathbf{F}_{\text{IR}}\}_{Q_{\text{tot}}=0}$ as the following mapping,
@@
where $\overline{\alpha}$ is the complex conjugation of $\alpha$ and we droped the parity indeces for simplicity. However, this only implies the condition for the charge and parity, and more detailed data are necessary for pratical calculations. Hence, the following discussion, we assume that the construction of the RG domain wall has been established by more respective techniques as in [102, 120, 104, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 15] and we concentrate our attention on the properties obtained from the charge and parity structures. Surprisingly, only from these structures, one can obtain nontrivial domain walls changing the chirality of anyons.
@@
The right arrow $\Rightarrow$ is obtained by bulk semionization or sandwich construction. To check consistency between this formalism and the domain wall Verlinde formula in [117] (or similar generalization [309]) is an interesting research direction. For simplicity, we study the models with anomalous UV and IR theories, but the same phenomenology should hold for the models with the anomaly free UV and IR theories.
@@
where $(i)$ is the label of two systems and $\pm$ is labeling chiral and anithiral components. Hence, by applying the folding trick[118], the two theories are Witt equivalent when $J_{\text{tot}}=J^{(1)}\overline{J}^{(2)}$ satisfy the anomaly-free condition.
@@
{ $\{\text{ Charged domain wall}\}\leftrightarrow\{\text{ Domain wall quark}\}$ (88)
@@ V Conclusion
Figure 10: Respective benefit of the bosonic representations and group extended representations. In a bosonic model, a correspondence between operators constructing correlation functions, quantum states, and (generalized) symmetry holds. Hence, one can achieve various calculations, by using vertex-operator-algebra or category theory and more. In a group-extended model analogous to a hadron model (or fractional supersymmetric model), the theory contains information on anomaly classifications determined by the conformal dimensions of the simple current. Moreover, one can attach gauge field easily. However, because the zero modes contain the nontrivial entanglement, calculations of the correlation functions become difficult.
@@
As a concluding remark, we note the respective benefits of the bosonic and fermionic (or group extended) representations. Bosonic models have been extensively studied because this representation usually provides systematic calculations of correlation functions and related physical observables. In a modern understanding, the correspondence between quantum states, correlators, and (generalized) symmetry seems to be a fundamental source of this benefit. On the other hand, the group-extended model corresponds to a particle (or quark-hadron) picture. Hence, one can attach gauge fields and apply the ideas of condensations more easily (Fig. 10). However, as we have observed in this work, the treatment of the zero modes and their entanglement structure requires careful treatment. We also remark on the difficulty of attaching gauge fields to bosonic systems[328, 329, 330, 331, 332, 333, 334, 335, 336, 337]. (It may be surprising for the readers, but the naive coupling of bosonic fields and $U(1)$ gauge field cannot reproduce the LSM theorem and the spectrum of twisted models[338, 339].) Because of the theoretical difficulties of the gauging operations, studies on the corresponding lattice or microscopic models are fundamentally important respectively. Inversely, the field-theoretical understanding of such models is also important for more general understanding of quantum phases.